There are two identical, positively charged conducting spheres fixed in space. The spheres are 39.4 cm apart (center to center) and repel each other with an electrostatic force of F1 = 0.0765 N. Then, a thin conducting wire connects the spheres, redistributing the charge on each sphere. When the wire is removed the spheres still repel but with a force of F2 = 0.100 N. Using this information, find the initial charge on each sphere, q1 and q2 if initially q1
Answer #1

here ,

let the charges are x and y

Now, for force needed

9 *10^9 * x * y/(0.394)^2 = 0.0765

Now, when they are touched

the charges will be (x + y)/2

9 *10^9 * ((x +y)/2)^2/.394^2 = 0.100

solving for x and y

x = 6.7 *10^-7 C

y = 1.95 *10^-6 C

the charges q1 and q2 are 6.7 *10-7 C and 1.95 *10-6 C